### v2

@@ Abstract
Understanding the structure of real data is paramount in advancing modern deep-learning methodologies. Natural data such as images are believed to be composed of features organised in a hierarchical and combinatorial manner, which neural networks capture during learning. Recent advancements show that diffusion models can generate high-quality images, hinting at their ability to capture this underlying structure. We study this phenomenon in a hierarchical generative model of data. We find that the backward diffusion process acting after a time $t$ is governed by a phase transition at some threshold time, where the probability of reconstructing high-level features, like the class of an image, suddenly drops. Instead, the reconstruction of low-level features, such as specific details of an image, evolves smoothly across the whole diffusion process. This result implies that at times beyond the transition, the class has changed but the generated sample may still be composed of low-level elements of the initial image. We validate these theoretical insights through numerical experiments on class-unconditional ImageNet diffusion models. Our analysis characterises the relationship between time and scale in diffusion models and puts forward generative models as powerful tools to model combinatorial data properties.
1 Introduction
Figure 1: Images generated by a denoising diffusion probabilistic model starting from the top-left image and inverting the dynamics at different times $t$. $T$ corresponds to the time scale when the forward diffusion process converges to an isotropic Gaussian distribution. At small $t$, the class of the generated image remains unchanged, with only alterations of low-level features, such as the eyes of the leopard. After a characteristic time $t$, the class undergoes a phase transition and changes. However, some low-level attributes of the original image are retained to compose the new image. For instance, the wolf is composed with eyes, nose and ears similar to those of the leopard, and the butterfly inherits its colours and black spots.
@@ 1 Introduction
Many theoretical studies in machine learning assume that the data distribution is Gaussian, which is ill-suited to study compositionality. Conversely, we consider hierarchical generative models of data, inspired by models of formal grammar and statistical physics, for which the ability of diffusion-based models to compose can be described exactly.
Diffusion models add noise to images as time increases and learn the reverse denoising process to generate new samples. If some finite amount of noise is added, and the process is then reversed, we predict that (i) for small noise, only low-level features of the image change, as already noticed empirically Ho et al. (2020), (ii) there exists a threshold noise at which the probability of remaining in the same class suddenly drops to near-random chance and (iii) surprisingly, beyond that point, low-level features may persist and compose the element of a new class. Remarkably, these results appear already evident in examples such as Figure 1 and can be systematically quantified by considering the change of internal representations in state-of-the-art CNNs, whereby a remarkable correspondence to our predictions is found.
Overall, our analysis reveals that diffusion models act at different hierarchical levels of the data at different time scales within the diffusion process. It also puts forward hierarchical generative models as a valuable theoretical tool to address several unanswered questions in machine learning.
@@ 1.1 Our contributions
On the theoretical side, we consider models belonging to the class of $L$-level context-free grammars where the structure of the latent variables is tree-like. We show that the denoising dynamics of diffusion models can be solved exactly for these models of data using belief propagation (BP). In particular, the analytical treatment of BP describes the evolution of the probability of latent variables for different levels of corruption noise and reveals a phase transition for a threshold noise value when $L$ is large. We show numerically that the dynamics of latent variables is reflected in the hidden representation of deep networks previously trained on a supervised classification task on these data.
On the empirical side, we test these insights by performing a systematic study of the denoising diffusion dynamics on ImageNet. Following Ho et al. (2020), we invert the noising process at some time $t$, leading to novel noiseless images. We then analyse how the representation of a state-of-the-art convolutional architecture changes between the initial and newly generated images as a function of both time $t$ and depth of the representation. Our results confirm the presence of a sharp transition in the class at a given time or noise level.
@@ 1.2 Related work
Most of the theoretical work on diffusion models considers simple models of data. Under mild assumptions on the data distribution, diffusion models exhibit a sample complexity that scales exponentially with the data dimension Block et al. (2020); Oko et al. (2023). This curse of dimensionality can be mitigated through stronger distributional assumptions, such as considering data lying within a low-dimensional latent subspace De Bortoli (2022); Chen et al. (2023); Yuan et al. (2023), Gaussian mixture models Biroli & Mézard (2023); Shah et al. (2023); Cui et al. (2023), graphical models Mei & Wu (2023), or data distributions that can be factorised across scales Kadkhodaie et al. (2023). For multimodal distributions such as Gaussian mixtures, the backward dynamics exhibits a cross-over time when it concentrates toward one of the modes Biroli & Mézard (2023); Ambrogioni (2023); Raya & Ambrogioni (2024). Closer to our work, Okawa et al. (2023) consider synthetic compositional data to empirically show how diffusion models learn to generalise by composing different concepts. In contrast, we study data that are not only compositional but also hierarchically structured and make quantitative predictions on how diffusion models compose features at different scales.
Hierachical models of data structure
Generative models of data have a long history in linguistics Rozenberg & Salomaa (1997) and have more recently received attention in the context of machine learning theory. In supervised learning, deep networks can represent hierarchical tasks more efficiently than shallow networks Poggio et al. (2017) and can efficiently learn them from an information theory viewpoint Schmidt-Hieber (2020). For hierarchical models of data, correlations between the input data and the task are critical for learning Mossel (2016); Shalev-Shwartz et al. (2017); Malach & Shalev-Shwartz (2018; 2020) and the representations learnt by neural networks with gradient descent reflect the hidden latent variables of such models both in CNNs Cagnetta et al. (2023) and transformers Allen-Zhu & Li (2023). In this work, we use these hierarchical generative models of data to study the denoising dynamics of diffusion models theoretically.
@@ 2 Diffusion models and feature hierarchies
Figure 2: Left panel. Examples of images generated by reverting the diffusion process at different times $t$. Starting from the left images $x_{0}$ at time $t=0$, we generate samples $\hat{x}_{0}(t)\sim p_{\theta}(\hat{x}_{0}|x_{t})$ by first running the diffusion process up to time $t$ and then reverting it, as described in Section 2.2. At time $t=T$, $x_{T}$ corresponds to isotropic Gaussian noise and the generated image $\hat{x}_{0}(T)$ is uncorrelated from $x_{0}$. At intermediate times, instead, a sudden change of the image class is observed, while some lower-level features are retained. Right panel. Cosine similarity between the post-activations of the hidden layers of a ConvNeXt Base Liu et al. (2022) for the initial images $x_{0}$ and the synthesised ones $\hat{x}_{0}(t)$. Around $t\approx T/2$, the similarity between logits exhibits a sharp drop, indicating the change in class, while the hidden representations of the first layers change more smoothly. This indicates that certain low-level features from the original images are retained for composing the sampled images also after the class transition. To compute the cosine similarity, all activations are standardised, i.e., centred around the mean and scaled by the standard deviation computed on the 50000 images of the ImageNet-1k validation set. At each time, the values of the cosine similarity correspond to the maximum of their empirical distribution over $10000$ images ($10$ per class of ImageNet-1k).
This section introduces denoising diffusion probabilistic models and demonstrates how class-unconditional ImageNet diffusion models operate on image features across different hierarchical levels at different time scales.  [^1] The code for running the experiments on ImageNet is available at github.com/pcsl-epfl/forward-backward-diffusion.
@@ 2.1 Background on denoising diffusion models
Denoising diffusion probabilistic models (DDPMs) Ho et al. (2020) are generative models designed to sample from a distribution by reversing a step-by-step noise addition process. In particular, let $q(\cdot)$ represent the data distribution, and let $x_{0}$ be a sample drawn from this distribution, i.e., $x_{0}\sim q(x_{0})$. First, DDPMs consist of a forward process which is a Markov chain generating a sequence of noised data {$x_{t}\}_{1\leq t\leq T}$ by introducing isotropic Gaussian noise at each time step $t$ with a variance schedule $\beta_{t}\}_{1\leq t\leq T}$ as follows:
@@
with $\overline{\alpha_{t}}=\prod_{t^{\prime}=1}^{t}(1-\beta_{t^{\prime}})$ and $\eta\sim\mathcal{N}(0,\mathbb{I})$. By selecting the noise schedule such that $\overline{\alpha}_{t}\to 0$ as $t\to T$, the distribution of $x_{T}$ becomes an isotropic Gaussian distribution. Subsequently, DDPMs reverse this process by gradually removing noise in a backward process. In this process, the models learn Gaussian transition kernels $q(x_{t-1}|x_{t})$ by parametrising their mean and variance using a neural network with parameters $\theta$ as follows:
@@
After training, the learnt $p_{\theta}$ can be used to generate novel examples by initiating the process with $x_{T}\sim\mathcal{N}(0,\mathbb{I})$ and running it in reverse to obtain a sample from $q$. We refer the reader to Ho et al. (2020); Nichol & Dhariwal (2021); Dhariwal & Nichol (2021) for more details regarding the formulation of DDPMs and the technical aspects of the reverse transition kernels parameterisation with neural networks.
@@ 2.2 Forward-backward experiments
Previous studies on DDPMs noted that inverting the diffusion process at different times $t$ starting from an image $x_{0}$ results in samples $\hat{x}_{0}(t)\sim p_{\theta}(\hat{x}_{0}|x_{t})$ with distinct characteristics depending on the choice of $t$. Specifically, when conditioning on the noisy samples $x_{t}$’s obtained by diffusing images from the CelebA dataset, one finds that for small values of $t$, only fine details change Ho et al. (2020). We conduct a similar experiment using a class-unconditional DDPM introduced by Dhariwal & Nichol (2021) on the ImageNet dataset with 256x256 resolution.
@@
    Similarly to the findings in Ho et al. (2020), at small inversion times $t$, only local features change. Furthermore, the class of the sampled images remains consistent with that of the corresponding starting images, i.e., ${{\rm class}(\hat{x}_{0}(t))=\rm class}(x_{0})$ with high probability.
@@
Our theory, presented in Section 4 and 5, predicts how features at different hierarchical levels vary at different time scales of the diffusion dynamics in accordance with observations (i), (ii), and (iii).
@@ 2.3 ImageNet hidden representations
To quantify the qualitative observations mentioned earlier, we design an experiment using the empirically known fact that deep learning models learn hierarchical representations of the data, with complexity increasing as the architecture’s depth grows. This phenomenon holds true in both real Olah et al. (2020); LeCun et al. (2015); Zeiler & Fergus (2014) and synthetic scenarios Cagnetta et al. (2023); Allen-Zhu & Li (2020). Therefore, we use these internal representations as a proxy for the compositional structure of the data. We investigate how the hidden representations of a deep ConvNeXt Base model Liu et al. (2022), achieving 96.9% top-5 accuracy on ImageNet, change as a function of the inversion time $t$ and depth $\ell$ of the representation. In the right panel of Figure 2, we illustrate the value of the cosine similarity between the post-activations of every hidden layer of the ConvNeXt for the initial and generated images. We observe that:
@@
    The representations of early layers of the network, corresponding to low-level and localised features of the images, are the first to change at short diffusion times and evolve smoothly.
@@
To study the robustness of our results with respect to the architecture choice, in Appendix D, we report the same measurements using ResNet architectures with varying width and depth He et al. (2016). We find the same qualitative behaviour as the ConvNeXt in Figure 2.
We now present our theory, which predicts these observations. In Section 3, we introduce a generative model of data which mimics the hierarchical and compositional structure of images while being analytically tractable. In Section 4, we solve exactly the denoising process for this model with different levels of noising. Finally, in Section 5, we prove the existence of a phase transition at a critical noise value.
@@ 3 Hierarchical generative model of data
Figure 3: Sketch of the generative hierarchical model we study in the paper. In this example, depth $L=3$ and branching factor $s=2$. Different values of the input and latent variables are represented with different colours.
To model the combinatorial and hierarchical nature of images, we consider hierarchical generative models Mossel (2016); Shalev-Shwartz et al. (2017); Malach & Shalev-Shwartz (2018; 2020); DeGiuli (2019); Allen-Zhu & Li (2020); Cagnetta et al. (2023). Specifically, we study an $L$-level context-free grammar. It consists of a set of class labels $\mathcal{C}\equiv\left\{1,\dots,v\right\}$ and an alphabet $\mathcal{A}\,{\equiv}\,\left\{a_{1},\dots,a_{v}\right\}$ of $v$ features. Once the class label $\gamma$ is picked uniformly at random from $\mathcal{C}$, the data is generated iteratively from a set of production rules with branching factor $s$ at each layer $\ell$ (see  Figure 3, for an illustration):
@@
We consider the case of the Random Hierarchy Model (RHM) Cagnetta et al. (2023), for which the $m$ production rules of any latent variable or class are sampled uniformly at random among the $v^{s}$ possible ones without replacement. In this case, the total number of possible data produced per class is $m\cdot m^{s}\cdots m^{s^{L-1}}=m^{\frac{d-1}{s-1}}$, where the dimension $d$ is defined as $d=s^{L}$. In the following, we use the notation $X^{(\ell)}_{i}$ to indicate the variable at layer $\ell$ and position $i\in\{1,\dots,s^{L-\ell}\}$.
@@
4 Exact denoising of the RHM with message passing

### v3

@@ Abstract
Understanding the structure of real data is paramount in advancing modern deep-learning methodologies. Natural data such as images are believed to be composed of features organized in a hierarchical and combinatorial manner, which neural networks capture during learning. Recent advancements show that diffusion models can generate high-quality images, hinting at their ability to capture this underlying compositional structure. We study this phenomenon in a hierarchical generative model of data. We find that the backward diffusion process acting after a time $t$ is governed by a phase transition at some threshold time, where the probability of reconstructing high-level features, like the class of an image, suddenly drops. Instead, the reconstruction of low-level features, such as specific details of an image, evolves smoothly across the whole diffusion process. This result implies that at times beyond the transition, the class has changed, but the generated sample may still be composed of low-level elements of the initial image. We validate these theoretical insights through numerical experiments on class-unconditional ImageNet diffusion models. Our analysis characterizes the relationship between time and scale in diffusion models and puts forward generative models as powerful tools to model combinatorial data properties.
keywords:
diffusion models |||| data structure |||| compositionality |||| deep learning
4]4
@@
[…]
Diffusion models add noise to images as time increases and learn the reverse denoising process to generate new samples. In particular, if some finite amount of noise is added to an image and the process is then reversed, we observe that: (i) for small noise, only low-level features of the image change ; (ii) at a threshold noise, the probability of remaining in the same class suddenly drops to near-random chance; (iii) beyond that point, low-level features may persist and compose the element of a new class. While observation (i) is intuitive and was first noticed in (12), the fact that at large noise only low-level features may remain unchanged is surprising. We will show below that this property is expected for hierarchical data. These results appear already evident in examples such as Figure 1, and we systematically quantify them by considering the change of internal representations in state-of-the-art convolutional neural networks.
We show that the observations (i), (ii), and (iii) can be theoretically explained through generative models of data with a hierarchical and compositional structure (10), inspired by models of formal grammars and statistical physics. We demonstrate that for these models the Bayes optimal denoising can be described exactly using belief propagation on tree-like graphs. Remarkably, our analysis predicts and explains both the phase transition in the class (observation (ii)) and how lower-level features compose to generate new data before and after this transition (observations (i) and (iii)).
Overall, our results reveal that diffusion models act at different hierarchical levels of the data at different time scales within the diffusion process and establish hierarchical generative models as valuable theoretical tools to address several unanswered questions in machine learning.
Figure 1: Illustration of forward-backward experiments. Images generated by a denoising diffusion probabilistic model starting from the top-left image and inverting the dynamics at different times $t$. $T$ corresponds to the time scale when the forward diffusion process converges to an isotropic Gaussian distribution. At small $t$, the class of the generated image remains unchanged, with only alterations of low-level features, such as the eyes of the leopard. After a characteristic time $t$, the class undergoes a phase transition and changes. However, some low-level attributes of the original image are retained to compose the new image. For instance, the wolf is composed of eyes, nose, and ears similar to those of the leopard, and the butterfly inherits its colors and black spots.
@@ 0.1 Our contributions
We perform a systematic study of the denoising diffusion dynamics on ImageNet. We invert the noising process at some time $t$, leading to novel noiseless images. We then analyze how the representation of state-of-the-art convolutional architectures changes between the initial and newly generated images as a function of both time $t$ and depth of the representation. This analysis reveals the presence of a sharp transition in the class at a given time or noise level. Importantly, at times beyond the transition, when the class has changed, we find that the generated images may still be composed of low-level features of the original image.
To model theoretically the compositional structure of images, we consider hierarchical generative models of data where the structure of the latent variables is tree-like. We use belief propagation to study the optimal denoising dynamics for such data and obtain the evolution of latent variables’ probabilities for different levels of corruption noise. In the limit of a large tree depth, this analysis reveals a phase transition for the probability of reconstructing the root node of the tree – which represents the class label of a data point – at a specific noise threshold. Conversely, the probability of reconstructing low-level latent variables evolves smoothly throughout the denoising diffusion process. Thus, after the transition, low-level features of the original datum may persist in composing a generated element of a new class, as we empirically observe in ImageNet. Finally, we show numerically that the dynamics of the latent variables is reflected in the hidden representation of deep networks previously trained on a supervised classification task on these data.
Organization of the paper
In Section 1, we introduce denoising diffusion probabilistic models and present our large-scale experiments on ImageNet data. In Section 2, we define the hierarchical generative model of data that we study theoretically. In Section 3, we study the optimal denoising for these data using message-passing techniques and show that our model captures the experimental observations on real data. In Section 4, we perform a mean-field analysis of the optimal denoising process, obtaining an analytical prediction for the phase transition of the class at a critical noise value and for the reconstruction probabilities of lower-level features.
@@ 0.2 Related work
Forward-backward protocol in diffusion-based models
(12) introduced the “forward-backward” protocol to probe diffusion-based models, whereby an image with a controlled level of noise is then denoised using a reverse-time diffusion process. It led to the observation that “when the noise is small, all but fine details are preserved, and when it is large, only large scale features are preserved”. Although our work agrees with the first part of the statement, it disagrees with the second. Our work also provides a systematic quantification of the effects of forward-backward experiments, going beyond qualitative observations based on individual images as in (12). Specifically, we introduce quantitative observables that characterize changes in the latent features of images and perform extensive experiments with state-of-the-art models, averaging results over $10^{5}$ ImageNet samples. Such quantification is key to connecting with theory. The forward-backward protocol was also studied in (17) to speed up the generation process of images.
@@
Most of the theoretical work on diffusion models considers simple models of data. Under mild assumptions on the data distribution, diffusion models exhibit a sample complexity that scales exponentially with the data dimension (18, 19). This curse of dimensionality can be mitigated through stronger distributional assumptions, such as considering data lying within a low-dimensional latent subspace (20, 21, 22), Gaussian mixture models (23, 24, 25), graphical models (26), or data distributions that can be factorized across scales (27). For multimodal distributions such as Gaussian mixtures, the backward dynamics exhibits a cross-over time when it concentrates toward one of the modes (23, 28, 29). This cross-over is similar to our observation (ii) above if these modes are interpreted as classes. As demonstrated in SI Appendix 5, such models of data cannot reproduce our salient predictions and observations. Closer to our work, (30) considers synthetic compositional data to empirically show how diffusion models learn to generalize by composing different concepts. In contrast, we study data that are not only compositional but also hierarchically structured and make quantitative predictions on how diffusion models compose features at different scales.
Hierarchical models of natural data
Generative models of data have a long history of describing the structure of language and image data. In linguistics, formal grammars describe the syntactic structure of a language through a hierarchical tree graph (31). Similar ideas have been explored to decompose visual scenes hierarchically into objects, parts, and primitives (32) and have been formalized in pattern theory (33). These hierarchical models led to practical algorithms for semantic segmentation and scene understanding, as illustrated in, e.g., (34, 35, 36). Recent works propose a hierarchical decomposition of images, in which latent variables are wavelet coefficients at different scales (37, 27). In this case, the graph is not tree-like (27) – a conclusion that could stem from the specific choice of latent variables.
Hierachical models in machine learning theory
More recently, generative models of data received attention in the context of machine learning theory. In supervised learning, deep networks can represent hierarchical tasks more efficiently than shallow networks (5) and can efficiently learn them from an information theory viewpoint (7). For hierarchical models of data, correlations between the input data and the task are critical for learning (4, 38, 6, 39) and the representations learned by neural networks with gradient descent reflect the hidden latent variables of such models both in Convolutional Neural Networks (CNNs) (10) and transformers (40). In this work, we use these hierarchical generative models of data to study the denoising dynamics of diffusion models theoretically.
@@ 1 Diffusion models and feature hierarchies
Figure 2: Left panel. Examples of images generated by reverting the diffusion process at different times $t$. Starting from the left images $x_{0}$ at time $t=0$, we generate samples $\hat{x}_{0}(t)\sim p_{\theta}(\hat{x}_{0}|x_{t})$ by first running the diffusion process up to time $t$ and then reverting it, as described in Section 1.2. At time $t=T$, $x_{T}$ corresponds to isotropic Gaussian noise and the generated image $\hat{x}_{0}(T)$ is uncorrelated from $x_{0}$. At intermediate times, instead, a sudden change of the image class is observed, while some lower-level features are retained. Right panel. Cosine similarity between the post-activations of the hidden layers of a ConvNeXt Base (41) for the initial images $x_{0}$ and the synthesized ones $\hat{x}_{0}(t)$. Around $t\approx T/2$, the similarity between logits exhibits a sharp drop, indicating the change in class, while the hidden representations of the first layers change more smoothly. This indicates that certain low-level features from the original images are retained for composing the sampled images also after the class transition. To compute the cosine similarity, all activations are standardized, i.e., centered around the mean and scaled by the standard deviation computed on the 50000 images of the ImageNet-1k validation set. At each time, the values of the cosine similarity correspond to the maximum of their empirical distribution over $10000$ images ($10$ per class of ImageNet-1k).
This section introduces denoising diffusion probabilistic models and demonstrates how class-unconditional ImageNet diffusion models operate on image features across different hierarchical levels at different time scales. [^1] The code for running the experiments on ImageNet is available at github.com/pcsl-epfl/forward-backward-diffusion (42).
@@ 1.1 Background on denoising diffusion models
Denoising diffusion probabilistic models (DDPMs) (12) are generative models designed to sample from a distribution by reversing a step-by-step noise addition process. In particular, let $q(\cdot)$ represent the data distribution, and let $x_{0}$ be a sample drawn from this distribution, i.e., $x_{0}\sim q(x_{0})$. First, DDPMs consist of a forward process which is a Markov chain generating a sequence of noised data {$x_{t}\}_{1\leq t\leq T}$ by introducing isotropic Gaussian noise at each time step $t$ with a variance schedule $\beta_{t}\}_{1\leq t\leq T}$ as follows:
@@
with $\overline{\alpha_{t}}=\prod_{t^{\prime}=1}^{t}(1-\beta_{t^{\prime}})$ and $\eta\sim\mathcal{N}(0,\mathbb{I})$. By selecting the noise schedule such that $\overline{\alpha}_{t}\to 0$ as $t\to T$, the distribution of $x_{T}$ becomes an isotropic Gaussian distribution. Subsequently, DDPMs reverse this process by gradually removing noise in a backward process. In this process, the models learn Gaussian transition kernels $q(x_{t-1}|x_{t})$ by parametrizing their mean and variance using a neural network with parameters $\theta$ as follows:
@@
After training, the learned $p_{\theta}$ can be used to generate novel examples by initiating the process with $x_{T}\sim\mathcal{N}(0,\mathbb{I})$ and running it in reverse to obtain a sample from $q$. We refer the reader to (12, 43, 44) for more details regarding the formulation of DDPMs and the technical aspects of the reverse transition kernels parameterization with neural networks.
@@ 1.2 Forward-backward experiments
Previous studies on DDPMs noted that inverting the diffusion process at different times $t$ starting from an image $x_{0}$ results in samples $\hat{x}_{0}(t)\sim p_{\theta}(\hat{x}_{0}|x_{t})$ with distinct characteristics depending on the choice of $t$. Specifically, when conditioning on the noisy samples $x_{t}$’s obtained by diffusing images from the CelebA dataset, one finds that for small values of $t$, only fine details change (12). We conduct a similar experiment using a class-unconditional DDPM introduced by (44), on the ImageNet dataset with 256x256 resolution.
@@
    Similarly to the findings in (12), at small inversion times $t$, only local features change. Furthermore, the class of the sampled images remains consistent with that of the corresponding starting images, i.e., ${{\rm class}(\hat{x}_{0}(t))=\rm class}(x_{0})$ with high probability.
@@
Our theory, presented in Section 3 and 4, predicts how features at different hierarchical levels vary at different time scales of the diffusion dynamics in accordance with observations (i), (ii), and (iii).
@@ 1.3 ImageNet hidden representations
To quantify the qualitative observations mentioned earlier, we design an experiment using the empirically known fact that deep learning models learn hierarchical representations of the data, with complexity increasing as the architecture’s depth grows. This phenomenon holds true in both real (45, 46, 47) and synthetic scenarios (10, 48). Therefore, we use these internal representations as a proxy for the compositional structure of the data. We investigate how the hidden representations of a deep ConvNeXt Base model (41), achieving 96.9% top-5 accuracy on ImageNet, change as a function of the inversion time $t$ and depth $\ell$ of the representation. In the right panel of Figure 2, we illustrate the value of the cosine similarity between the post-activations of every hidden layer of the ConvNeXt for the initial and generated images. We observe that:
@@
    The representations of early layers of the network, corresponding to low-level and localized features of the images, are the first to change at short diffusion times and evolve smoothly.
@@
To study the robustness of our results with respect to the architecture choice, in SI Appendix 4, we report the same measurements using ResNet architectures with varying width and depth (49). We find the same qualitative behavior as the ConvNeXt in Figure 2.
We now present our theory, which predicts these observations.
@@ 2 Hierarchical generative model of data
input variables(level 00)low latents(level $1$)high latents(level $2$)label
Figure 3: Sketch of the hierarchical and compositional structure of data. Left panel: The leopard in the image can be iteratively decomposed in features at different levels of abstraction. Right panel: Generative hierarchical model we study in this paper. In this example, depth $L=3$ and branching factor $s=2$. Different values of the input and latent variables are represented with different colors.
In this section, we introduce a generative model of data that mimics the structure of images while being analytically tractable. Natural images often display a hierarchical and compositional structure (50). Take, for example, the image of a snow leopard (see Figure 3). This image is composed of multiple high-level components, such as the head and the paws. Each of these components, in turn, is composed of sub-features. For instance, the head comprises elements like ears, eyes, and mouth. Further dissecting these elements, we find even more granular details, such as edges that define the finer aspects of each feature. To model this hierarchical and compositional nature of images, we consider hierarchical generative models (4, 38, 6, 39, 51, 48, 10). In particular, consider a set of class labels $\mathcal{C}\equiv\left\{1,\dots,v\right\}$ and an alphabet $\mathcal{A}\,{\equiv}\,\left\{a_{1},\dots,a_{v}\right\}$ of $v$ features. Once the class label $\gamma$ is picked uniformly at random from $\mathcal{C}$, the data is generated iteratively from a set of production rules with branching factor $s$ at each layer $\ell$ (see  Figure 3, for an illustration):
@@
We consider the case of the Random Hierarchy Model (RHM) (10), for which the $m$ production rules of any latent variable or class are sampled uniformly at random among the $v^{s}$ possible ones without replacement. In this case, the total number of possible data produced per class is $m\cdot m^{s}\cdots m^{s^{L-1}}=m^{\frac{d-1}{s-1}}$, which has exponential dependence in the dimension $d=s^{L}$. In the following, we use the notation $X^{(\ell)}_{i}$ to indicate the variable at layer $\ell$ and position $i\in\{1,\dots,s^{L-\ell}\}$.
@@
3 Optimal denoising of the RHM with message passing